### v1

# PGOT: A Physics-Geometry Operator Transformer for Complex PDEs

Abstract
While Transformers have demonstrated remarkable potential in modeling Partial Differential Equations (PDEs), modeling large-scale unstructured meshes with complex geometries remains a significant challenge. Existing efficient architectures often employ feature dimensionality reduction strategies, which inadvertently induces Geometric Aliasing, resulting in the loss of critical physical boundary information. To address this, we propose the Physics-Geometry Operator Transformer (PGOT), designed to reconstruct physical feature learning through explicit geometry awareness. Specifically, we propose Spectrum-Preserving Geometric Attention (SpecGeo-Attention). Utilizing a “physics slicing-geometry injection” mechanism, this module incorporates multi-scale geometric encodings to explicitly preserve multi-scale geometric features while maintaining linear computational complexity $O(N)$. Furthermore, PGOT dynamically routes computations to low-order linear paths for smooth regions and high-order non-linear paths for shock waves and discontinuities based on spatial coordinates, enabling spatially adaptive and high-precision physical field modeling. PGOT achieves consistent state-of-the-art performance across four standard benchmarks and excels in large-scale industrial tasks including airfoil and car designs.
Machine Learning, ICML
1 Introduction
Partial Differential Equations (PDEs), derived from fundamental principles such as conservation of mass and energy, are ubiquitous in modeling complex phenomena across science and engineering, including fluid dynamics, structural mechanics, and climate modeling (Cao et al., 2025). Traditional numerical methods, such as the Finite Element Method (FEM) (Zienkiewicz & Taylor, 2005) and Finite Volume Method (FVM) (Eymard et al., 2000), approximate solutions by discretizing PDEs onto meshes (Chen et al., 2025). However, their reliance on dense mesh discretization imposes a heavy computational burden, often requiring days or even months for multi-physics problems necessitating fine spatiotemporal resolution. Recently, Neural PDE Solvers have emerged as a promising paradigm, capable of learning resolution-invariant solution mappings between infinite-dimensional function spaces (Wang et al., 2025a; Li et al., 2025). By training on offline data, these operators can predict families of solutions under varying parameters orders of magnitude faster than traditional solvers, offering a promising direction for real-time scientific computing.
Figure 1: Efficiency and accuracy comparison on standard benchmarks. (a) Inference speed vs. Memory usage. Bubble size indicates model size. (b) Multi-dimensional performance metrics on PDE and industrial datasets.
Early spectral-based neural operators, such as the Fourier Neural Operator (FNO) (Li et al., 2021), rely on the fast fourier transform. While they possess the remarkable ability to learn resolution-independent operators, they are inherently limited to regular grids, restricting their applicability to complex geometries and unstructured meshes. To address this, graph-based neural operators (Hao et al., 2023; Li et al., 2023b) treat input and output functions as graphs, employing message-passing mechanisms to generalize across different domain geometries. However, graph methods are often constrained by local receptive fields, making it difficult to effectively capture global long-range dependencies and suffering from low efficiency on large-scale nodes.
This necessitates the adoption of Transformers. The self-attention mechanism in Transformers is mathematically analogous to learning a global Green’s function (Cao, 2021), making it ideally suited for modeling non-local physical interactions (Alkin et al., 2024; Koshizuka et al., 2024). However, a fundamental scalability barrier remains: the standard self-attention mechanism exhibits quadratic computational complexity $\mathcal{O}(N^{2})$ with respect to the number of mesh points $N$. Given that scientific simulations typically involve $10^{5}$ to $10^{7}$ grid points, standard Transformers are computationally intractable for high-resolution physics. Consequently, the research frontier has shifted towards efficient linear transformers (Li et al., 2023c; Luo et al., 2025; Wu et al., 2024). Recent state-of-the-art methods employ strategies such as low-rank approximations or token clustering to reduce the effective sequence length, achieving linear complexity $\mathcal{O}(N)$ (Rao et al., 2021; Xiao et al., 2024). While these architectures successfully improve throughput, they often do so at the cost of geometric and physical fidelity, as follows:
L1): Aggregating local mesh points into latent tokens implicitly acts as a spatial low-pass filter, causing geometric aliasing where high-frequency boundary details are erased.
L2): Allocating a uniform computational budget across the domain neglects physical heterogeneity, resulting in over-smoothed shocks in discontinuity regions and redundant computations in smooth flows.
To bridge the gap between computational efficiency and high-fidelity physical modeling, we present the Physics-Geometry Operator Transformer (PGOT). Unlike prior methods that treat geometry merely as a supplementary input, PGOT intertwines physical state aggregation with explicit geometric reconstruction. Specifically, PGOT first projects physical fields into compact latent slices, while explicitly injecting multi-scale geometric embeddings during the reconstruction phase. This eliminates geometric aliasing caused by conventional dimensionality reduction while maintaining $\mathcal{O}(N)$ linear complexity (Solving L1). Subsequently, the model dynamically routes computational paths based on local flow complexity. This design ensures the model is both spectrum-preserving to avoid aliasing and spatially adaptive to handle physical heterogeneity (Solving L2). Extensive experiments demonstrate that PGOT achieves state-of-the-art performance on standard benchmarks and industrial applications, as illustrated by the efficiency-accuracy trade-off in Figure 1.
Our main contributions are as follows:
  • Our proposed PGOT framework addresses the geometric aliasing bottleneck inherent in complex mesh modeling, thereby reconciling computational efficiency with physical fidelity.
  • We propose the SpecGeo-Attention mechanism to decouple global physical aggregation from local geometric reconstruction via a “physics slicing-geometry injection” paradigm, thereby preserving complex boundary features with high fidelity.
  • By dynamically routing computations to linear or non-linear paths based on flow heterogeneity, the Taylor-decomposed Feed-Forward Network efficiently captures multi-scale physical features while enhancing model interpretability.
2 Related Work
2.1 Physics-Informed Neural Networks
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) represent a foundational paradigm for solving PDEs by embedding governing equations directly into the loss function. PINNs parameterize the solution as a neural network and minimize the PDE residual at collocation points, enabling mesh-free training without labeled simulation data. However, PINNs face fundamental limitations for the problems we target. First, PINNs solve individual PDE instances, requiring complete retraining for each new parameter configuration or boundary condition—a prohibitive cost for design optimization requiring thousands of evaluations. Second, for complex geometries and multi-physics coupling, PINNs frequently encounter ill-conditioning issues (Zhang et al., 2025), often leading to convergence failures or accuracy degradation. Although recent research has alleviated these problems through parameterization (Cho et al., 2024), meta-learning (Cheng & Alkhalifah, 2025), and transfer learning (Wang et al., 2025b) strategies, PINNs still struggle to maintain accuracy and efficiency when the configuration space becomes sufficiently large. These limitations motivate the development of the neural operator paradigm.
2.2 Neural Operators for PDEs
Neural operators have emerged as a powerful paradigm for learning resolution-invariant mappings between infinite-dimensional function spaces (Kovachki et al., 2023). DeepONet (Lu et al., 2021) pioneered this field by approximating operators through a Branch-Trunk architecture. Subsequently, the Fourier Neural Operator (FNO) (Li et al., 2021) established new benchmarks by parameterizing the integral kernel directly in the frequency domain using FFT. To address the limitations of standard FNO, numerous variants have been proposed: U-NO (Rahman et al., 2023) and U-FNO (Wen et al., 2022) integrate U-Net structures to capture multi-scale features, while F-FNO (Tran et al., 2023) improves computational efficiency by factorizing the Fourier representation across dimensions. To handle irregular geometries beyond rectangular grids, Geo-FNO (Li et al., 2023a) maps physical domains to latent uniform grids, while LSM (Wu et al., 2023) and LNO (Wang & Wang, 2024) apply spectral methods within learned low-dimensional latent spaces to tackle high-dimensional PDE problems. Furthermore, to mitigate the global nature of Fourier transforms that typically leads to loss of local details, wavelet-based methods such as WMT (Gupta et al., 2021) and WNO (Tripura & Chakraborty, 2023) have been introduced to leverage the time-frequency localization properties of wavelets for capturing local variations. For modeling complex unstructured meshes, geometric deep learning offers an alternative perspective. Graph Neural Operators (GNO) (Li et al., 2020) and MeshGraphNet (Pfaff et al., 2021) treat physical domains as graphs, employing message-passing mechanisms to handle complex boundaries. To combine global spectral processing capabilities with local geometric flexibility, hybrid architectures have emerged. For instance, GINO (Li et al., 2023b) integrates Geo-FNO’s global spectral capabilities with GNO’s local graph processing, while 3D-GeoCA (Deng et al., 2024) further incorporates pre-trained 3D vision backbones to enhance geometric feature extraction. However, graph-based methods typically suffer from geometric instability (Morris et al., 2023; Luo et al., 2025) and incur high computational overhead when scaling to million-scale mesh points due to neighborhood search and message passing.
2.3 Transformers for PDEs
The Transformer architecture has garnered significant attention in scientific computing due to its global receptive field and superior modeling capabilities. Early works such as OFormer (Li et al., 2023c) and HT-Net (Liu et al., 2024) demonstrated the potential of attention mechanisms in capturing non-local physical interactions. However, the quadratic complexity $\mathcal{O}(N^{2})$ of standard attention remains a bottleneck restricting its scalability. To achieve linear complexity $\mathcal{O}(N)$, recent research has explored diverse pathways: AFNO (Guibas et al., 2022), FNet  (Lee-Thorp et al., 2022), and GFNet (Rao et al., 2021) attempt token mixing in the Fourier domain, while SAOT (Zhou et al., 2025) further integrates wavelet attention to mitigate aliasing issues; simultaneously, Galerkin Transformer (Cao, 2021), GNOT (Hao et al., 2023), and ONO (Xiao et al., 2024) utilize kernel tricks or low-rank approximations to achieve linearization; whereas methods represented by FactFormer  (Li et al., 2023d) and Transolver (Wu et al., 2024) employ axial decomposition or token aggregation strategies, respectively. Despite their significant efficiency improvements, these dimensionality reduction or aggregation operations act as implicit spatial low-pass filters, inevitably discarding high-frequency geometric details. Furthermore, they typically allocate a uniform computational budget across the entire domain, failing to adapt to physical heterogeneity in flow fields.
3 Method
In this section, we present the design of the Physics-Geometry Operator Transformer (PGOT). We first analyze the geometric aliasing bottleneck inherent in existing efficient architectures. Subsequently, we detail the two core components: Spectrum-Preserving Geometric Attention (SpecGeo-Attention) and the Taylor-Decomposed Feed-Forward Network (TaylorDecomp-FFN). The overall architecture is illustrated in Figure 2.
Figure 2: Overall architecture of PGOT. (a) The framework explicitly integrates multi-scale geometry via stacked PhysGeoBlocks to reconstruct velocity and pressure fields on complex 3D meshes. (b) Visualization of TaylorDecomp-FFN. The Linear Expert (blue) captures smooth conservation dynamics, while the Non-linear Expert (red) targets high-order fluctuations.
Figure 3: Architecture of PhysGeoBlock and SpecGeo-Attention. (a) The PhysGeoBlock integrates explicit geometric coordinates into the SpecGeo-Attention and TaylorDecomp-FFN layers. (b) The “physics slicing-geometry injection” paradigm. A Spectrum Encoder generates geometry-aware weights to aggregate NN mesh points into $M$ latent tokens (Slice) and reconstruct them (DeSlice). This design achieves linear complexity $\mathcal{O}(N)$ while preserving multi-scale geometric fidelity.
3.1 Problem Formulation
We consider a physical system governed by partial differential equations (PDEs), where the solution operator $\mathcal{G}^{\dagger}$ maps an input function space to an output function space: $\mathcal{G}^{\dagger}:\mathcal{A}(\Omega;\mathbb{R}^{d_{a}})\rightarrow\mathcal{U}(\Omega;\mathbb{R}^{d_{u}})$. In practice, the continuous domain $\Omega\subset\mathbb{R}^{d}$ is discretized into $N$ mesh points $\mathbf{G}=\{\mathbf{g}_{i}\}_{i=1}^{N}$. The input and output functions are sampled as $\mathbf{a}\in\mathbb{R}^{N\times d_{a}}$ and $\mathbf{u}\in\mathbb{R}^{N\times d_{u}}$, respectively. Our goal is to learn a parameterized neural operator $\mathcal{G}_{\theta}$ that approximates the ground truth operator $\mathcal{G}^{\dagger}$.
Existing efficient Transformers achieve linear complexity through feature aggregation. However, this operation acts as a low-pass filter in the frequency domain. Let $f(\mathbf{g})$ be the geometric signal. Aggregation can be modeled as a convolution with a smoothing kernel $K_{\sigma}$. According to the convolution theorem, the frequency response is $\hat{f}_{agg}(\omega)=\hat{K}_{\sigma}(\omega)\cdot\hat{f}(\omega)$. When the spatial frequency $|\omega|$ exceeds the cutoff frequency $\omega_{c}\propto 1/\sigma$, $\hat{K}_{\sigma}(\omega)\approx 0$, causing high-frequency geometric details to be irreversibly erased. We term this phenomenon geometric aliasing, identifying it as a fundamental bottleneck limiting performance on complex boundaries. PGOT addresses this via explicit multi-scale geometric reconstruction.
3.2 PGOT
Recent work has demonstrated that slice-based attention, which aggregates mesh points into learnable physics-aware tokens, can achieve linear complexity for PDE solving (Wu et al., 2024). However, this aggregation implicitly acts as a spatial low-pass filter, causing geometric aliasing at complex boundaries. To address this limitation, PGOT introduces Spectrum-Preserving Geometric Attention and Taylor-Decomposed Feed-Forward Networks.
$\mathbf{X}^{0}=\mathcal{P}_{\theta}\left(\mathbf{a}\oplus\mathcal{E}_{\text{pos}}(\mathbf{G})\right),$ (1)
where $\oplus$ denotes concatenation, and $\mathcal{P}_{\theta}$ is a lifting MLP. We employ a Unified Reference Point Encoding $\mathcal{E}_{\text{pos}}$ to ensure discretization invariance.
The processor consists of $L$ stacked PhysGeoBlocks, as depicted in Figure 3(a). Unlike prior methods that use coordinates only at the input, PGOT explicitly injects geometry $\mathbf{G}$ at every layer:
$\hat{\mathbf{X}}^{l}$ $=\mathbf{X}^{l-1}+\text{SpecGeo-Attn}\left(\text{LN}(\mathbf{X}^{l-1});\mathbf{G}\right),$ (2)
$\mathbf{X}^{l}$ $=\hat{\mathbf{X}}^{l}+\text{TaylorDecomp}\left(\text{LN}(\hat{\mathbf{X}}^{l});\mathbf{G}\right),$ (3)
where $\text{LN}(\cdot)$ denotes Layer Normalization. This design ensures that geometric guidance persists throughout the deep network evolution.
3.2.1 Spectrum-Preserving Geometric Attention
As illustrated in Figure 3(b), the core philosophy of SpecGeo-Attention is to decouple global physical state aggregation from local geometric reconstruction.
Multi-Scale Geometric Spectral Encoding.
To reconstruct high-frequency information lost during aggregation, we inject multi-scale geometric encodings prior to slicing. We design a bank of encoders for $S$ scales:
$\mathbf{h}_{s}=\phi_{s}\left(10^{s-1}\cdot\mathbf{g}\right),\quad s=1,\ldots,S,$ (4)
where $\phi_{s}$ is a two-layer MLP. The exponential scaling factor $10^{s-1}$ ensures coverage of diverse spatial frequencies: $s=1$ captures macro-structures, while higher scales capture micro-boundary details. These features are fused via concatenation to preserve full spectral information:
$\Phi_{\text{geo}}(\mathbf{g})=\sigma_{\text{act}}\left(\mathbf{W}_{\text{fuse}}\cdot\left[\mathbf{h}_{1};\mathbf{h}_{2};\ldots;\mathbf{h}_{S}\right]\right),$ (5)
where $\cdot$$\cdot$] denotes concatenation.
Geometry-Informed Physics Slicing.
The geometric encoding is injected into the physical features to guide the slicing process:
$\tilde{\mathbf{X}}_{q}=\mathbf{X}\mathbf{W}_{x}+\Phi_{\text{geo}}(\mathbf{G}).$ (6)
We then compute the soft assignment matrix $\mathbf{A}\in\mathbb{R}^{N\times M}$ mapping $N$ mesh points to $M$ latent physical states:
$\mathbf{A}_{ij}=\frac{\exp\left([\tilde{\mathbf{X}}_{q}]_{i}\cdot\mathbf{w}_{j}/\tau\right)}{\sum_{k=1}^{M}\exp\left([\tilde{\mathbf{X}}_{q}]_{i}\cdot\mathbf{w}_{k}/\tau\right)},$ (7)
where $\{\mathbf{w}_{j}\}$ are learnable slice prototypes. Since $\mathbf{A}$ is derived from geometry-enhanced features, the clustering respects complex boundaries (e.g., distinguishing the pressure side from the suction side of an airfoil).
Physical features are then aggregated into latent tokens $\mathbf{Z}\in\mathbb{R}^{M\times C}$ via $\mathbf{Z}=\mathbf{D}^{-1}\mathbf{A}^{\top}(\mathbf{X}\mathbf{W}_{f})$, processed by standard Multi-Head Self-Attention (MHSA), and reconstructed back to the mesh domain:
$\mathbf{X}^{\prime}=\mathbf{A}\cdot\text{MHSA}(\mathbf{Z}).$ (8)
This entire process maintains $\mathcal{O}(N)$ complexity while preserving geometric fidelity.
3.2.2 TaylorDecomp-FFN
Physical fields exhibit significant spatial heterogeneity: smooth laminar regions follow low-order linear laws, while shocks and discontinuities are governed by high-order non-linear dynamics. Drawing inspiration from Taylor series expansion, we approximate the solution operator at any spatial point $\mathbf{g}$ by decomposing it into low-order and high-order components:
$\mathbf{u}(\mathbf{g})\approx\underbrace{\mathbf{u}(\mathbf{g}_{0})+\nabla\mathbf{u}|_{\mathbf{g}_{0}}(\mathbf{g}-\mathbf{g}_{0})}_{\text{Linear Terms}}+\underbrace{\mathcal{R}_{high}(\mathbf{g})}_{\text{Non-linear Residuals}}.$ (9)
To adaptively model these terms, we propose the TaylorDecomp-FFN, which formulates the feature transformation as a dynamic composition of two expert paths:
$\mathbf{X}_{out}=(\mathbf{1}-\boldsymbol{\alpha}(\mathbf{g}))\odot\mathcal{F}_{\text{lin}}(\mathbf{X})+\boldsymbol{\alpha}(\mathbf{g})\odot\mathcal{F}_{\text{non}}(\mathbf{X}).$ (10)
The architecture explicitly instantiates the theoretical decomposition above. The Linear Expert ($\mathcal{F}_{\text{lin}}$), corresponding to the low-order Taylor terms, consists of a streamlined linear projection with dropout to efficiently capture smooth conservation dynamics:
$\mathcal{F}_{\text{lin}}(\mathbf{x})=\mathbf{W}_{l2}(\text{Dropout}(\mathbf{W}_{l1}\mathbf{x})).$ (11)
In parallel, the Non-linear Expert ($\mathcal{F}_{\text{non}}$) targets the high-order residuals, utilizing a deeper MLP with activation functions to resolve sharp gradients and discontinuities:
$\mathcal{F}_{\text{non}}(\mathbf{x})=\mathbf{W}_{n2}(\sigma(\mathbf{W}_{n1}\mathbf{x})).$ (12)
The dynamic routing between these paths is governed by a spatial gate $\boldsymbol{\alpha}(\mathbf{g})\in[0,1]^{C}$, which is generated via an Absolute Coordinate Embedding (ACE):
$\boldsymbol{\alpha}(\mathbf{g})=\text{Sigmoid}\left(\text{MLP}_{gate}(\text{PosEmbed}(\mathbf{g}))\right).$ (13)
We employ unconstrained regime selection where the gate $\boldsymbol{\alpha}$ operates independently per channel. This flexibility allows the model to fully activate the non-linear path in shock regions ($\alpha\to 1$), rely on the linear path in smooth flows ($\alpha\to 0$), or leverage a hybrid combination ($\alpha\approx 0.5$) in complex transition zones.
4 Experiments
We evaluate PGOT on four standard benchmarks and two large-scale industrial simulations, covering diverse geometric settings from point clouds to complex 3D unstructured meshes.
4.1 Experimental Setup
Datasets.
Our evaluation encompasses six benchmarks with varying geometric complexities. The standard benchmarks (Li et al., 2021, 2023a) include Elasticity (inner stress prediction on 2D point clouds), Plasticity (temporal displacement forecasting under external forces), Airfoil (transonic flow simulation with shock waves), and Pipe (velocity field estimation in curved structures). For industrial-scale validation, we employ Shape-Net Car (Umetani & Bickel, 2018) for automotive aerodynamic simulation on hybrid 3D geometries, and AirfRANS (Bonnet et al., 2022) for high-fidelity RANS simulation on NACA airfoils under varying Reynolds numbers and angles of attack. Detailed dataset descriptions are provided in Appendix A.1.
Baselines.
We compare against representative methods from three categories: spectral neural operators (FNO (Li et al., 2021), Geo-FNO (Li et al., 2023a), U-NO (Rahman et al., 2023), F-FNO (Tran et al., 2023), LSM (Wu et al., 2023), WMT (Gupta et al., 2021)), Transformer-based approaches (Galerkin Transformer (Cao, 2021), OFormer (Li et al., 2023c), GNOT (Hao et al., 2023), ONO (Xiao et al., 2024), FactFormer (Li et al., 2023d), IPOT (Lee & Oh, 2024), Transolver (Wu et al., 2024), SAOT (Zhou et al., 2025)), and geometric deep learning methods (PointNet (Qi et al., 2017), Graph U-Net (Gao & Ji, 2019), GraphSAGE (Hamilton et al., 2017), MeshGraphNet (Pfaff et al., 2021), GNO (Li et al., 2020), GINO (Li et al., 2023b), 3D-GeoCA (Deng et al., 2024)).
Metrics and Implementation.
We adopt the relative $L^{2}$ error as the primary evaluation metric. For industrial benchmarks, we additionally report aerodynamic coefficient errors and Spearman’s rank correlation. PGOT is configured with $L=8$ PhysGeoBlocks, with other hyperparameters adapted to each task. All experiments are conducted on NVIDIA A100 GPUs and repeated three times. Comprehensive details on metrics, training configurations, and hyperparameters are provided in Appendix A.2.
4.2 Main Results
Table 1: Performance comparison on standard benchmarks. The relative $L^{2}$ error is reported. The best results are highlighted in bold, and the second best are underlined. Gray background indicates our proposed method. Improvement denotes the reduction in relative error.
Model Structured Mesh Point Cloud
Plasticity Pipe Airfoil Elasticity
WMT (2021) 0.0076 0.0077 0.0075 0.0359
U-FNO (2022) 0.0039 0.0056 0.0269 0.0239
Geo-FNO (2023a) 0.0074 0.0067 0.0138 0.0229
U-NO (2023) 0.0034 0.0100 0.0078 0.0258
F-FNO (2023) 0.0047 0.0070 0.0078 0.0263
LSM (2023) 0.0025 0.0050 0.0059 0.0218
Galerkin (2021) 0.0120 0.0098 0.0118 0.0240
HT-Net (2024) 0.0333 0.0059 0.0065 /
OFormer (2023c) 0.0017 0.0168 0.0183 0.0183
GNOT (2023) 0.0336 0.0047 0.0076 0.0086
FactFormer (2023d) 0.0312 0.0060 0.0071 /
ONO (2024) 0.0048 0.0052 0.0061 0.0118
IPOT (2024) 0.0033 / 0.0088 0.0156
Transolver (2024) 0.0013 0.0043 0.0053 0.0079
SAOT (2025) 0.0014 0.0062 0.0051 0.0090
PGOT (Ours) 0.0012 0.0039 0.0046 0.0069
Improvement 7.7% 9.3% 9.8% 12.7%
Table 2: Performance comparison on practical design tasks. For Shape-Net Car, relative $L^{2}$ of surrounding (Volume) and surface (Surf) physics fields is recorded. For AirfRANS, MSE of Volume and Surf fields is recorded. The relative $L^{2}$ of drag coefficient ($C_{D}$) and lift coefficient ($C_{L}$) is also recorded, along with their Spearman’s rank correlations $\rho_{D}$ and $\rho_{L}$. A Spearman’s correlation close to 1 indicates better performance. “/” indicates the baseline cannot apply to this benchmark.
Model Shape-Net Car AirfRANS
Volume $\downarrow$ Surf $\downarrow$ $C_{D}$ $\downarrow$ $\rho_{D}$ $\uparrow$ Volume $\downarrow$ Surf $\downarrow$ $C_{L}$ $\downarrow$ $\rho_{L}$ $\uparrow$
Simple MLP 0.0512 0.1304 0.0307 0.9496 0.0081 0.0200 0.2108 0.9932
GraphSAGE (2017) 0.0461 0.1050 0.0270 0.9695 0.0087 0.0184 0.1476 0.9964
PointNet (2017) 0.0494 0.1104 0.0298 0.9583 0.0253 0.0996 0.1973 0.9919
Graph U-Net (2019) 0.0471 0.1102 0.0226 0.9725 0.0076 0.0144 0.1677 0.9949
MeshGraphNet (2021) 0.0354 0.0781 0.0168 0.9840 0.0214 0.0387 0.2252 0.9945
GNO (2020) 0.0383 0.0815 0.0172 0.9834 0.0269 0.0405 0.2016 0.9938
Galerkin (2021) 0.0339 0.0878 0.0179 0.9764 0.0074 0.0159 0.2336 0.9951
Geo-FNO (2023a) 0.1670 0.2378 0.0664 0.8280 0.0361 0.0301 0.6161 0.9257
GNOT (2023) 0.0329 0.0798 0.0178 0.9833 0.0049 0.0152 0.1992 0.9942
GINO (2023b) 0.0386 0.0810 0.0184 0.9826 0.0297 0.0482 0.1821 0.9958
3D-GeoCA (2024) 0.0319 0.0779 0.0159 0.9842 / / / /
Transolver (2024) 0.0231 0.0787 0.0113 0.9901 0.0037 0.0142 0.1030 0.9978
SAOT (2025) 0.0223 0.0778 0.0122 0.9884 0.0038 0.0128 0.0872 0.9983
PGOT (Ours) 0.0210 0.0766 0.0101 0.9926 0.0025 0.0024 0.0766 0.9990
Improvement 9.1% 1.5% 10.6% / 32.4% 81.3% 12.2% /
Standard Benchmarks.
Table 1 presents the performance comparison on four standard benchmarks. PGOT consistently achieves state-of-the-art results across all datasets, with relative error reductions ranging from 7.7% to 12.7% compared to the second-best methods. On Plasticity and Pipe, PGOT outperforms Transolver by 7.7% and 9.3%, respectively, demonstrating the effectiveness of explicit geometric injection in capturing temporal dynamics and curved boundary features. For Airfoil, where shock waves introduce sharp discontinuities, PGOT achieves 9.8% improvement over SAOT, validating the benefit of TaylorDecomp-FFN in adaptively routing computations to non-linear paths for discontinuity regions. This is qualitatively supported by Figure 4, where PGOT produces substantially cleaner predictions near shock waves compared to Transolver and SAOT. The most significant improvement (12.7%) is observed on Elasticity, a point cloud benchmark where geometric information is particularly sparse. As visualized in Figure 4 (right), baselines exhibit scattered high-error points, whereas PGOT maintains consistently low errors, highlighting the importance of multi-scale geometric encoding in preserving structural details.
Figure 4: Visualization of prediction errors on standard benchmarks. The first column shows ground truth fields, and the remaining columns display prediction errors.
Industrial Simulations.
Table 2 summarizes results on large-scale industrial benchmarks. On Shape-Net Car, PGOT achieves the best performance across all metrics, with 9.1% reduction in volumetric field error and 10.6% improvement in drag coefficient prediction compared to Transolver. The Spearman correlation $\rho_{D}=0.9926$ indicates excellent ranking capability for design optimization. Figure 5(b) visualizes these results, showing excellent agreement between PGOT’s predicted streamlines and ground truth, with superior accuracy in geometrically complex regions like wheel wells. On AirfRANS, PGOT demonstrates substantial improvements: 32.4% reduction in volumetric field error, 81.3% reduction in surface field error, and 12.2% improvement in lift coefficient prediction. As shown in Figure 5(a), while Transolver and SAOT produce radial error patterns, PGOT achieves significantly cleaner predictions with errors primarily concentrated in the far-field region. These results underscore PGOT’s capability in handling complex unstructured meshes with over 32,000 points while maintaining high prediction fidelity.
Figure 5: Visualization on industrial benchmarks. (a) AirfRANS: ground truth pressure field and prediction errors. (b) Shape-Net Car: ground truth streamlines and PGOT prediction, along with surrounding velocity and surface pressure errors.
4.3 Ablation Studies
We conduct systematic ablation studies to validate the contribution of each proposed component. We report results on four standard benchmarks (Plasticity, Pipe, Airfoil, Elasticity) in the main text to cover diverse geometric types and physical regimes; full results covering all datasets are provided in Appendix B.1.
Table 3: Ablation on the number of latent slices $M$. We observe that $M=64$ offers the best trade-off for standard benchmarks.
Slices $M$ Plasticity Pipe Airfoil Elasticity
8 0.0104 0.0079 0.0194 0.0239
16 0.0047 0.0071 0.0101 0.0074
32 0.0025 0.0040 0.0049 0.0081
64 0.0012 0.0039 0.0046 0.0069
128 0.0021 0.0051 0.0056 0.0061
256 0.0134 0.0044 0.0048 0.0079
Number of Latent Slices $M$.
Table 3 examines the impact of varying $M$. Performance generally improves as $M$ increases, with Plasticity, Pipe, and Airfoil peaking at $M=64$. Interestingly, Elasticity benefits from even finer granularity, reaching its best performance at $M=128$ (0.0061). This is likely because point cloud data lacks explicit connectivity, requiring more slices to capture local geometric nuances. However, further increasing $M$ to 256 leads to degradation across all tasks due to the fragmentation of coherent physical regions. Considering the trade-off between accuracy and computational cost, we adopt $M=64$ as the unified default.
Table 4: Ablation on the number of geometric scales $S$. Multi-scale encoding ($S=2$) significantly outperforms single-scale baselines.
Scales $S$ Plasticity Pipe Airfoil Elasticity
1 0.0104 0.0079 0.0194 0.0239
2 0.0012 0.0039 0.0046 0.0069
3 0.0047 0.0071 0.0101 0.0074
4 0.0021 0.0051 0.0056 0.0079
Number of Geometric Scales $S$.
Table 4 investigates the effect of multi-scale geometric encoding. Single-scale encoding ($S=1$) performs poorly across all benchmarks, confirming that aggregation-induced geometric aliasing significantly degrades prediction quality. Performance improves substantially and reaches optimality at $S=2$, demonstrating that a dual-scale representation is sufficient to capture both global topology and local boundary details. Interestingly, further increasing the scales to $S=3$ or $S=4$ leads to performance degradation. This suggests that excessive high-frequency geometric priors may introduce spectral noise or redundancy, which interferes with the model’s ability to learn robust physical dynamics.
Table 5: Both SpecGeo-Attention (SGA) and TaylorDecomp-FFN (TDF) are critical for optimal performance.
Variant Plasticity Pipe Airfoil Elasticity
w/o SGA 0.0093 0.0045 0.0104 0.0074
w/o TDF 0.0134 0.0089 0.0094 0.0231
Full PGOT 0.0012 0.0039 0.0046 0.0069
Component Analysis.
Table 5 validates the necessity of both proposed modules. Removing SpecGeo-Attention (w/o SGA) leads to performance drops across all tasks, particularly in Elasticity ($0.0069\to 0.0074$) and Airfoil ($0.0046\to 0.0104$), confirming that explicit geometric injection is crucial for handling irregular point clouds and shock boundaries. Removing TaylorDecomp-FFN (w/o TDF) causes severe degradation in Plasticity ($0.0012\to 0.0134$), validating the importance of adaptive routing for time-dependent heterogeneous fields.
4.4 Out-of-Distribution (OOD) Generalization
We evaluate PGOT on AirfRANS under two extrapolation settings: (1) Unseen Reynolds and (2) Unseen Angles. As shown in Table 6, PGOT demonstrates superior robustness against 11 baselines. In Reynolds extrapolation, PGOT achieves the lowest lift coefficient error ($C_{L}=0.2942$) and the highest ranking correlation ($\rho_{L}=0.9923$). In Angle extrapolation, PGOT maintains the highest ranking consistency ($\rho_{L}=0.9952$), significantly outperforming graph-based methods and surpassing the SOTA Transolver. This confirms that explicitly modeling geometry-physics interactions enables robust reasoning in unseen physical regimes. Detailed analysis are provided in Appendix B.3.
Table 6: OOD generalization on AirfRANS.
Model Unseen Reynolds Unseen Angles
$C_{L}$ $\downarrow$ $\rho_{L}$ $\uparrow$ $C_{L}$ $\downarrow$ $\rho_{L}\uparrow$
Simple MLP 0.6205 0.9578 0.4128 0.9572
GraphSAGE (2017) 0.4333 0.9707 0.2538 0.9894
PointNet (2017) 0.3836 0.9806 0.4425 0.9784
Graph U-Net (2019) 0.4664 0.9645 0.3756 0.9816
MeshGraphNet (2021) 1.7718 0.7631 0.6525 0.8927
GNO (2020) 0.4408 0.9878 0.3038 0.9884
Galerkin (2021) 0.4615 0.9826 0.3814 0.9821
GNOT (2023) 0.3268 0.9865 0.3497 0.9868
GINO (2023b) 0.4180 0.9645 0.2583 0.9923
Transolver (2024) 0.3042 0.9899 0.1503 0.9948
SAOT (2025) 0.4242 0.9764 0.2925 0.9888
PGOT (Ours) 0.2942 0.9923 0.1742 0.9952
5 Conclusion
In this work, we proposed PGOT to address the geometric aliasing bottleneck in efficient neural operators. By synergizing Spectrum-Preserving Geometric Attention with a regime-adaptive Taylor-Decomposed FFN, our framework reconciles linear complexity with high-fidelity boundary reconstruction. PGOT achieves state-of-the-art results on industrial benchmarks and demonstrates superior robustness in extrapolation tasks. Future avenues include extending this geometric-physical decoupling paradigm to multi-physics foundation models.
Appendix A Implementation Details
A.1 Dataset Descriptions
We provide detailed descriptions of all six benchmarks used in our evaluation. Table 7 summarizes the key statistics.
Table 7: Summary of benchmark datasets. #Mesh denotes the number of discretized mesh points per sample.
Geometry Benchmark Dim #Mesh Input Output Train / Test
Point Cloud Elasticity 2D 972 Structure Inner Stress 1000 / 200
Structured Mesh Plasticity 2D+T 3,131 External Force Displacement 900 / 80
Airfoil 2D 11,271 Structure Mach Number 1000 / 200
Pipe 2D 16,641 Structure Velocity 1000 / 200
Unstructured Mesh Shape-Net Car 3D 32,186 Structure Velocity, Pressure 789 / 100
AirfRANS 2D 32,000 Structure Velocity, Pressure 800 / 200
Elasticity.
This benchmark involves predicting the inner stress distribution of elastic materials based on geometric structures, discretized as 972 scattered points in 2D space. Each sample represents a different material structure, and the model must estimate the stress tensor at each point. The training set contains 1000 samples, and the test set contains 200 samples (Li et al., 2021; Wu et al., 2024).
Plasticity.
The task is to predict future deformations of plastic materials under impact from arbitrarily-shaped dies. The input consists of structured meshes with $101\times 31$ points representing the die geometry, and the output is the displacement field over 20 future time steps containing four directional deformation components. The training set contains 900 samples, and the test set contains 80 samples (Li et al., 2021; Wu et al., 2024).
Airfoil.
This benchmark requires estimating Mach number distributions around airfoil geometries on structured meshes with $221\times 51$ points. All shapes are deformations of the NACA-0012 baseline airfoil. The task involves transonic flow regimes where shock waves introduce discontinuities in the solution field. The training set contains 1000 samples, and the test set contains 200 samples (Li et al., 2021; Wu et al., 2024).
Pipe.
The objective is to predict horizontal fluid velocity fields based on curved pipe structures. Each sample is discretized into a $129\times 129$ structured mesh, with different samples generated by varying the pipe centerline geometry. The training set contains 1000 samples, and the test set contains 200 samples (Li et al., 2021; Wu et al., 2024).
Shape-Net Car.
This industrial benchmark focuses on automotive aerodynamic simulation at 72 km/h driving conditions. The task requires simultaneous prediction of surface pressure and surrounding air velocity. Vehicle geometries are sourced from the “car” category of the ShapeNet repository and discretized into unstructured meshes with approximately 32,000 points combining surface and volumetric elements. The input is preprocessed to include mesh point positions, signed distance functions, and normal vectors. The predicted fields can be used to compute drag coefficients for design optimization. The training set contains 789 samples, and the test set contains 100 samples (Umetani & Bickel, 2018).
AirfRANS.
This dataset contains high-fidelity Reynolds-Averaged Navier-Stokes simulation data for airfoils from the NACA 4- and 5-digit series. Unlike the Airfoil benchmark, AirfRANS features diverse airfoil shapes under varying Reynolds numbers ($3\times 10^{6}$ to $6\times 10^{6}$) and angles of attack ($5\textdegree$ to $15\textdegree$), with finer unstructured meshes of 32,000 points per sample. Air velocity, pressure, and viscosity are recorded for surrounding regions, along with surface pressure. Both lift and drag coefficients can be computed from the predicted fields. However, as noted in the original paper, air velocity estimation remains challenging for aircraft, causing all deep models to perform poorly on drag coefficient prediction. Therefore, we focus primarily on lift coefficient and volume/surface physical quantity prediction accuracy. The training set contains 800 samples, and the test set contains 200 samples (Bonnet et al., 2022).
A.2 Evaluation Metrics
Relative $L^{2}$ Error.
Given the ground-truth physical field $\mathbf{u}$ and the model prediction $\hat{\mathbf{u}}$, the relative $L^{2}$ error is defined as:
$\text{Relative }L^{2}=\frac{\|\mathbf{u}-\hat{\mathbf{u}}\|_{2}}{\|\mathbf{u}\|_{2}}.$ (14)
This metric is used across all benchmarks to measure the accuracy of predicted physical fields.
Aerodynamic Coefficient Error.
For Shape-Net Car and AirfRANS, we compute aerodynamic coefficients from the predicted fields and report their relative errors. For unit density fluid, the coefficient (drag or lift) is defined as:
$C=\frac{2}{v^{2}A}\left[\int_{\partial\Omega}p(\xi)\left(\hat{\mathbf{n}}(\xi)\cdot\hat{\mathbf{i}}(\xi)\right)d\xi+\int_{\partial\Omega}\tau(\xi)\cdot\hat{\mathbf{i}}(\xi)d\xi\right],$ (15)
where $v$ is the inlet flow speed, $A$ is the reference area, $\partial\Omega$ denotes the object surface, $p$ is the pressure function, $\hat{\mathbf{n}}$ is the outward unit normal vector of the surface, $\hat{\mathbf{i}}$ is the inlet flow direction, and $\tau$ represents the wall shear stress on the surface. The wall shear stress $\tau$ can be calculated from the air velocity near the surface, which is typically much smaller than the pressure term. For the drag coefficient of Shape-Net Car, $\hat{\mathbf{i}}=(-1,0,0)$ and $A$ is the area of the smallest rectangle enclosing the front of the car. For the lift coefficient of AirfRANS, $\hat{\mathbf{i}}=(0,0,-1)$. The relative $L^{2}$ error is computed between the ground-truth coefficient and the coefficient calculated from the predicted fields.
Spearman’s Rank Correlation.
Given $K$ test samples with ground-truth coefficients $C=\{C^{1},\cdots,C^{K}\}$ (drag or lift) and model-predicted coefficients $\hat{C}=\{\hat{C}^{1},\cdots,\hat{C}^{K}\}$, Spearman’s correlation coefficient is defined as the Pearson correlation coefficient between the rank variables:
$\rho=\frac{\text{cov}(R(C),R(\hat{C}))}{\sigma_{R(C)}\sigma_{R(\hat{C})}}$, (16)
where $R$ is the ranking function, cov denotes the covariance, and $\sigma$ represents the standard deviation of the rank variables. This metric measures the correlation between the ranking distribution of ground-truth coefficients and model-predicted values across all test samples, quantifying the model’s ability to correctly rank different designs. This property is particularly important for shape optimization applications. A Spearman’s correlation close to 1 indicates better performance.
A.3 Training and Model Configurations
Following the conventions established in prior work (Li et al., 2021, 2023a; Wu et al., 2024; Zhou et al., 2025), our model is trained using the relative $L^{2}$ loss. Table 8 summarizes the training configurations shared across all methods for fair comparison, and Table 9 presents PGOT-specific hyperparameters.
Table 8: Training configurations for each benchmark. All methods are trained under identical settings.
Benchmark Loss Epochs Learning Rate Optimizer Batch Size
Elasticity Relative $L^{2}$ 500 $10^{-3}$ AdamW 1
Plasticity Relative $L^{2}$ 500 $10^{-3}$ AdamW 8
Airfoil Relative $L^{2}$ 500 $10^{-3}$ AdamW 4
Pipe Relative $L^{2}$ 500 $10^{-3}$ AdamW 4
Shape-Net Car $\mathcal{L}_{v}+0.5\mathcal{L}_{s}$ 200 $10^{-3}$ Adam 1
AirfRANS $\mathcal{L}_{v}+\mathcal{L}_{s}$ 400 $10^{-3}$ Adam 1
For Shape-Net Car (Umetani & Bickel, 2018) and AirfRANS (Bonnet et al., 2022), $\mathcal{L}_{v}$ and $\mathcal{L}_{s}$ denote the relative $L^{2}$ losses on volumetric and surface fields.
Table 9: PGOT model configurations for each benchmark.
Benchmark Layers $L$ Channels $C$ Slices $M$ Scales $S$ Heads
Elasticity 8 128 64 2 8
Plasticity 8 128 64 2 8
Airfoil 8 128 64 2 8
Pipe 8 128 64 2 8
Shape-Net Car 8 256 32 3 8
AirfRANS 8 256 32 3 8
The hidden dimension $C$ is set to 256 for high-dimensional output tasks (Shape-Net Car and AirfRANS) and 128 for others. Accordingly, the number of latent slices $M$ is adjusted to 32 for $C=256$ and 64 for $C=128$ to balance model capacity and computational efficiency. These settings are consistent with the configurations used in Transolver (Wu et al., 2024) and other neural operators, and their effectiveness is further validated by our ablation studies.
A.4 Baseline Implementations
Neural Operators.
For FNO (Li et al., 2021), Geo-FNO (Li et al., 2023a), WMT (Gupta et al., 2021), U-FNO (Wen et al., 2022), U-NO (Rahman et al., 2023), F-FNO (Tran et al., 2023), and LSM (Wu et al., 2023), we report results from the original publications (Wu et al., 2024; Zhou et al., 2025; Li et al., 2021).
Transformer-based Methods.
For Galerkin Transformer (Cao, 2021), OFormer (Li et al., 2023c), GNOT (Hao et al., 2023), HT-Net (Liu et al., 2024), FactFormer (Li et al., 2023d), and ONO, we report results from the original publications (Wu et al., 2024; Zhou et al., 2025; Li et al., 2021). For Transolver (Wu et al., 2024) and SAOT (Zhou et al., 2025), we reproduce results based on their official codebases with all hyperparameter configurations kept consistent with PGOT to ensure fair comparison.
Geometric Deep Learning Methods.
For MLP, GraphSAGE (Hamilton et al., 2017), PointNet (Qi et al., 2017), Graph U-Net (Gao & Ji, 2019), MeshGraphNet (Pfaff et al., 2021), GNO (Li et al., 2020), GINO (Li et al., 2023b), and 3D-GeoCA (Deng et al., 2024), we report results from the original publications.
Appendix B Supplementary Results
B.1 Ablation Results
We provide the complete ablation study results on all six benchmarks in Tables 10, 11, and 12.
Table 10: Full ablation on the number of latent slices $M$.
$M$ Plasticity Pipe Airfoil Elasticity Car (Vol.) AirfRANS (Vol.)
8 0.0104 0.0079 0.0194 0.0239 0.0421 0.0077
16 0.0047 0.0071 0.0101 0.0074 0.0394 0.0059
32 0.0025 0.0040 0.0060 0.0081 0.0210 0.0025
64 0.0012 0.0039 0.0046 0.0069 0.0299 0.0055
128 0.0021 0.0051 0.0056 0.0061 0.0322 0.0049
256 0.0134 0.0044 0.0048 0.0079 0.0318 0.0104
Table 11: Full ablation on the number of geometric scales $S$.
$S$ Plasticity Pipe Airfoil Elasticity Car (Vol.) AirfRANS (Vol.)
1 0.0104 0.0079 0.0194 0.0239 0.0342 0.0029
2 0.0012 0.0039 0.0046 0.0069 0.0237 0.0031
3 0.0047 0.0071 0.0101 0.0074 0.0210 0.0025
4 0.0021 0.0051 0.0056 0.0079 0.0329 0.0098
Table 12: Full ablation on key components.
Variant Plasticity Pipe Airfoil Elasticity Car (Vol.) AirfRANS (Vol.)
w/o SGA 0.0093 0.0045 0.0104 0.0074 0.0482 0.0081
w/o TDF 0.0134 0.0089 0.0094 0.0231 0.0348 0.0076
Full PGOT 0.0012 0.0039 0.0046 0.0069 0.0210 0.0025
B.2 Efficiency Analysis
[…]
Table 13: Efficiency comparison of Transformer-based methods across different mesh sizes NN. Running time is measured by the time (s) to complete one epoch containing 1000 iterations. Params denotes the number of model parameters (MB).
Model Params GPU Memory (GB) Time (s/epoch)
1024 2048 4096 8192 16384 32768 1024 2048 4096 8192 16384 32768
Galerkin (2021) 1.041 0.62 0.66 0.74 0.91 1.45 2.05 26.51 26.50 27.48 37.10 67.52 129.87
OFormer (2023c) 0.884 0.63 0.69 0.80 1.02 1.67 2.44 28.15 30.98 31.11 47.90 91.67 182.21
ONO (2024) 1.109 1.47 1.75 2.30 3.47 5.64 10.09 69.76 76.25 100.13 149.59 255.34 462.46
GNOT (2023) 5.248 0.85 1.07 1.47 2.33 4.23 7.46 54.28 55.94 60.86 67.17 112.55 209.92
Transolver (2024) 0.950 0.15 0.27 0.53 1.02 2.03 4.02 38.43 38.56 37.33 38.36 51.73 98.54
SAOT (2025) 1.492 0.33 0.52 0.92 1.61 3.11 6.33 94.09 94.72 97.01 97.22 132.33 234.94
PGOT (Ours) 1.148 0.20 0.37 0.71 1.38 2.71 5.37 58.04 59.76 59.52 60.44 60.17 110.99
B.3 Out-of-Distribution Generalization
We evaluate PGOT under two extrapolation settings on AirfRANS: (1) Unseen Reynolds numbers and (2) Unseen angles of attack. The detailed experimental settings are provided in Table 14, where the training and test sets contain completely different Reynolds numbers or angles of attack.
Table 14: Settings of OOD generalization experiments on AirfRANS. The range of Reynolds numbers and angles of attack are listed.
Dataset Unseen Reynolds Unseen Angles
Reynolds Range Samples Angles Range Samples
Training Set $3\times 10^{6},5\times 10^{6}]$ 500 $2.5\textdegree,12.5\textdegree]$ 800
Test Set $2\times 10^{6},3\times 10^{6}]\cup[5\times 10^{6},6\times 10^{6}]$ 500 $5\textdegree,-2.5\textdegree]\cup[12.5\textdegree,15\textdegree]$ 200
Table 15: Full OOD generalization results on AirfRANS. PGOT achieves the best overall performance, particularly in ranking consistency ($\rho_{L}$) and field reconstruction (Volume/Surf). Best results are bolded, second best are underlined.
Model Unseen Reynolds Numbers Unseen Angles of Attack
Volume $\downarrow$ Surf $\downarrow$ $C_{L}$ $\downarrow$ $\rho_{L}$ $\uparrow$ Volume $\downarrow$ Surf $\downarrow$ $C_{L}$ $\downarrow$ $\rho_{L}$ $\uparrow$
Simple MLP 0.0669 0.1153 0.6205 0.9578 0.1309 0.3311 0.4128 0.9572
GraphSAGE (2017) 0.0798 0.1254 0.4333 0.9707 0.1192 0.2359 0.2538 0.9894
PointNet (2017) 0.0838 0.1403 0.3836 0.9806 0.2021 0.4649 0.4425 0.9784
Graph U-Net (2019) 0.0538 0.1168 0.4664 0.9645 0.0979 0.2391 0.3756 0.9816
MeshGraphNet (2021) 0.2789 0.2382 1.7718 0.7631 0.4902 1.1071 0.6525 0.8927
GNO (2020) 0.0833 0.1562 0.4408 0.9878 0.1626 0.2359 0.3038 0.9884
Galerkin (2021) 0.0330 0.0972 0.4615 0.9826 0.0577 0.2773 0.3814 0.9821
GNOT (2023) 0.0305 0.0959 0.3268 0.9865 0.0471 0.3466 0.3497 0.9868
GINO (2023b) 0.0839 0.1825 0.4180 0.9645 0.1589 0.2469 0.2583 0.9923
Transolver (2024) 0.0234 0.0341 0.3042 0.9899 0.0454 0.2279 0.1503 0.9948
SAOT (2025) 0.1032 0.1193 0.4242 0.9764 0.0748 0.2744 0.2925 0.9888
PGOT (Ours) 0.0203 0.0293 0.2942 0.9923 0.0348 0.2203 0.1742 0.9952
Table 15 presents the comprehensive evaluation of 12 models on AirfRANS extrapolation tasks. Beyond the Lift Coefficient ($C_{L}$) and its ranking correlation ($\rho_{L}$) reported in the main text, we additionally provide the Mean Squared Error (MSE) for surrounding (Volume) and surface (Surf) physics fields to offer a holistic view of model robustness.
Performance Analysis. Graph-based methods (e.g., MeshGraphNet (Pfaff et al., 2021), Graph U-Net (Gao & Ji, 2019)) exhibit significant degradation in extrapolation settings. For instance, MeshGraphNet incurs a relative error of 1.7718 for $C_{L}$ in Reynolds extrapolation. Fundamentally, these Message Passing Neural Networks tend to overfit the local connectivity statistics of the training mesh. When the Reynolds number increases, the physical boundary layer thins, introducing sharp gradients that alter the effective physical length scales; Message Passing Neural Networks, constrained by learned local receptive fields, fail to adapt to this shift. While Transformer-based baselines like Transolver (Wu et al., 2024) and SAOT (Zhou et al., 2025) improve upon this by leveraging global attention, they still rely heavily on latent feature similarity for token clustering. In scenarios with severe flow separation, the feature distribution shifts drastically, leading to unstable slicing and degraded field reconstruction.
In contrast, PGOT achieves strong stability across most metrics, obtaining the best performance in 7 out of 8 evaluation criteria. While Transolver slightly outperforms PGOT on $C_{L}$ in the Unseen Angles setting (0.1503 vs. 0.1742), PGOT achieves superior ranking correlation ($\rho_{L}$: 0.9952 vs. 0.9948), which is more critical for design optimization as it measures the model’s ability to correctly rank different airfoil designs. Moreover, PGOT consistently achieves the lowest field reconstruction errors (Volume and Surf) in both OOD settings, demonstrating that its geometric anchoring mechanism provides more reliable physical field predictions even when the coefficient prediction is slightly less accurate.
Appendix C Additional Visualizations
C.1 Learned Slice Assignments of SpecGeo-Attention
Figures 611 visualize the learned slice assignments from SpecGeo-Attention across all six benchmarks. Each subplot represents one of 32 latent physical states, where different colors indicate varying assignment intensities.
Figure 6: Learned slice assignments from SpecGeo-Attention on Airfoil (32 slices).
Figure 7: Learned slice assignments from SpecGeo-Attention on Pipe (32 slices).
Figure 8: Learned slice assignments from SpecGeo-Attention on Plasticity (32 slices).
Figure 9: Learned slice assignments from SpecGeo-Attention on Elasticity (32 slices).
Figure 10: Learned slice assignments from SpecGeo-Attention on AirfRANS (32 slices).
Figure 11: Learned slice assignments from SpecGeo-Attention on Shape-Net Car (32 slices).
C.2 Gate Activations of TaylorDecomp-FFN
Figures 1217 visualize the gate activations $\boldsymbol{\alpha}(\mathbf{g})$ from TaylorDecomp-FFN. Each subplot displays the gate values for one of 64 feature channels. Red indicates high activation (routing to the non-linear expert) and blue indicates low activation (routing to the linear expert).
Figure 12: Gate activations from TaylorDecomp-FFN on Airfoil (64 channels).
Figure 13: Gate activations from TaylorDecomp-FFN on Pipe (64 channels).
Figure 14: Gate activations from TaylorDecomp-FFN on Plasticity (64 channels).
Figure 15: Gate activations from TaylorDecomp-FFN on Elasticity (64 channels).
Figure 16: Gate activations from TaylorDecomp-FFN on AirfRANS (64 channels).
Figure 17: Gate activations from TaylorDecomp-FFN on Shape-Net Car (64 channels).